OpenStudy (anonymous):

\left[ \begin{matrix} 4 & -2 \\ -7 & 2 \end{matrix}\right] X=\left[\begin{matrix} -6\\12 \end{matrix} \right]

OpenStudy (anonymous):

\[\left[ \begin{matrix} 4 & -2 \\ -7 & 2 \end{matrix}\right] X=\left[\begin{matrix} -6\\12 \end{matrix} \right]\]

OpenStudy (anonymous):

\[\left[\begin{matrix} 4& -2\\-7&2 \end{matrix}\right]\]

OpenStudy (anonymous):

close the \left[

OpenStudy (anonymous):

yep I forgot to do that earlier :S I know I have to find the inverse somehow...where do I start?

OpenStudy (anonymous):

oh i was just going to solve like a donky \[4x_1-2x_2=-6\]\[-7x_1+2x_2=12\]

OpenStudy (anonymous):

this one is pretty easy, add and get \[3x_1=-6\implies x_1=-2\] etc

OpenStudy (anonymous):

would we get -2 -1